OpenStudy (anonymous):

Help & show work: Find the length of the following segment created by these pairs of endpoints: (0, 1), (6, 6)

OpenStudy (angela210793):

\[d=\sqrt{(x2-x1)^2+(y2-y1)^2}\]

OpenStudy (anonymous):

over 6 (from zero to one) up 5 use pythagoras \[c^2=6^2+5^2=25+36=61\] so distance is \[\sqrt{61}\]

OpenStudy (angela210793):

just replace x1;y1 coordiaantes of the 1st point and x2;y2 coordinates of the 2nd point

OpenStudy (anonymous):

A: 7.8 B: 9.1 C: 2.5 D: 6

OpenStudy (angela210793):

7.8

OpenStudy (anonymous):

Well how do i solve this?

OpenStudy (angela210793):

i gave u the formula

OpenStudy (anonymous):

can you brake it down? a bit more plz

OpenStudy (amistre64):

the steps are: subtract one point from the other: (6, 6) -(0, 1) ------ 6 ,5 ; use these as the lengths of the legs of a right triangle, with the hypotenuse being the distance between the points

OpenStudy (amistre64):

d^2 = 6^2 + 5^2 d^2 = 36 + 25 d^2 = 61 ; to find d sqrt it all d = sqrt(61) , and simplifiy

OpenStudy (anonymous):

Umm you lost me at 6,5 :S

OpenStudy (anonymous):

maybe i need more sleep :(

OpenStudy (amistre64):

ok , what does 6 and 5 represent as far as distances between th epoints?

OpenStudy (amistre64):

do you see how I got 6 and 5?

OpenStudy (anonymous):

yes but once i get that what do i do simple math plz

OpenStudy (amistre64):

the simplest math possible is the pythagorean thrm for this ..

OpenStudy (anonymous):

o jod

OpenStudy (amistre64):

once you know the distance across the base (the difference in xs) and the distance up the side (the difference in ys) you have all the parts needed to use the pythag thrm

OpenStudy (anonymous):

so XS is 6 or the 5 in 6, 5

OpenStudy (anonymous):

and ys is the 6 or the 5 in 6,5

OpenStudy (amistre64):

at this point it doesnt matter does it? does: 6^2 + 5^2 = 5^2 + 6^2 ??

OpenStudy (amistre64):

in other words, does 36+25 = 25+36 ?? if so, then its mathematically moot

OpenStudy (anonymous):

:'( i'm lost sorry i think i need more sleep

OpenStudy (amistre64):

get some sleep and look at it again with fresh buggery eyes :)

OpenStudy (amistre64):

if you draw a picture it might help to place things

OpenStudy (amistre64):